# Just Functioning as a Hook for Two-Stage Referring Multi-Object Tracking

### v2

@@ 3.5 Loss Computation
We supervise all intermediate and final results using a multi-class cross-entropy loss similar to Focal Loss [36] to enhance the model’s multi-scale capability and alleviate the imbalanced sample distribution in the dataset:
$L=-\sum_{N}[\alpha\cdot y_{n}+(1-\alpha)(1-y_{n})]\cdot l_{h}$ (3)
where, $\alpha$ is used to balance recall and precision, while $l_{h}$ is used for hard sample mining, and its calculation is as follows:
$l_{h}=l_{ce}\cdot\texttt{detach}(h_{n}^{\beta}),\quad h_{n}$ $=-\sum_{c=1}^{C}p(O_{n,c})\log p(O_{n,c})$ (4)
where, $h_{n}$ is the information entropy of prediction and $\beta$ is used to control its strength. detach refers to the “detach" operation in PyTorch, which removes $h_{n}$ from the computation graph to prevent it from affecting the back-propagation.
4 Experiment
4.1 Experiment Setup
Dataset and Metrics
@@ 4.1 Experiment Setup
We report the performance of JustHook on the only two widely used datasets: Refer-KITTI [4] and Refer-KITTI-V2 [5] in the manuscript, and on Refer-Dance [8] in E.1. For Refer-KITTI, we use a small subset of labels from Refer-KITTI-V2 to correct the original annotations. We further discuss it in the Appendix. For a fair comparison, we use Higher Order Tracking Accuracy [37] (HOTA) as the main metric, following TransRMOT [4]. HOTA comprehensively considers the detection and association capabilities of the tracking model. We also report the localization accuracy (LocA), the detection accuracy (DetA) and the association accuracy (AssA).
Implement Details
@@
Benefited from the hook-like framework, JustHook can function as a plug-and-play model that connects with any visual backbone and text encoder. The visual backbones we used include ResNet [9], Swin [10; 38], and RoPE Swin [34], all of which use the pre-trained parameters on ImageNet [39], with an input size of $224\times 672$ to align as closely as possible with the aspect ratio of KITTI [40]. For text encoders, we test both RoBERTa [35] and BERT [33].
In each epoch, we iterate over all frames of all targets, with sample stride 2, and each sample consists of 4 frames. During inference, the sample stride is set to 1. In each training iteration, each target is coupled with 8 randomly sampled expressions, among which 4 ones are positive and 4 ones are negative. We use AdamW [41] as the optimizer, with a learning rate of 3e-5, and train for 20 epochs. But in many cases, JustHook only requires 10 epochs to achieve good performance. All experiments are performed on two NVIDIA 4090 GPUs.
4.2 Benchmark Experiments
We compare our method with SOTA approaches on Refer-KITTI and Refer-KITTI-V2 in Table 1. The trackers coupled with JustHook all follow the TBD paradigm, and we report only the results for the best detector & tracker here: TempRMOT [5] & NeuralSORT [8]. As a detector, TempRMOT is trained using detection labels, where each target in each frame is treated as a new target, and it only requires 10 epochs instead of 60 [5] to converge. Structurally, it can be seen as a DeformableDETR [42] with added inter-frame interactions. We further compare and discuss the impacts of different trackers [43; 44; 8; 42] on the performance of JustHook in C.
Compared to the current SOTA method TempRMOT, our best model, JustHook + RoBERTa [35] + ROPE Swin-T [34], achieves superior performance on both Refer-KITTI and Refer-KITTI-V2. JustHook has a greater advantage on Refer-KITTI-V2, surpassing TempRMOT by 7.77%, 8.38%, and 5.03% on HOTA, DetA, and AssA, respectively. JustHook particularly exhibits significant advantages in HOTA and DetA, demonstrating its powerful matching capability, especially on the complex Refer-KITTI-V2 dataset. We further report the performance of JustHook on Refer-Dance in E.1.
Compared with iKUN, it can be found that JustHook with the same CLIP encoder can achieve improvements of 4.09% and 31.32% on Refer-KITTI and Refer-KITTI-V2 respectively. Since in the semantic space of CLIP, the role of cosine similarity has more advantages compared with PCD, this demonstrates that VFH can make more efficient use of the knowledge of pre-trained models compared with reusing encoders and additional modules. Meanwhile, VFH adaptively extracts local features from the original backbone, causing less disruption to the original semantic space of CLIP and being more natural compared with iKUN, enabling VFH to achieve significant performance improvements on the complex Refer-KITTI-V2 dataset.
Qualitative Results
The qualitative results can be found in F.
4.3 Ablation Study
Optim-flow-based Motion Extractor
As shown in Table 3, it seems that more complex aggregation methods perform even worse than the simple MLP-based approach. This can be attributed to the fact that pre-trained visual backbones already encode rich static information within each frame. Complex aggregation strategies can distort the feature space, making subsequent learning more difficult.
Table 2: Comparison of the performance between various sequential decoding and PCD.
$Q_{1}$ $KV_{1}$ $Q_{2}$ $KV_{2}$ HOTA DetA AssA
[…]
Cosine Similarity 28.29 13.46 59.58
PCD 42.81 31.35 58.61
@@ 4.3 Ablation Study
Attention 41.38 29.76 57.69
Conv 41.90 29.79 59.06
MLP 42.81 31.35 58.61
Parallel Combined Decoder
In Table 3, we investigate the performance of PCD. Among them, "Cosine Similarity" means removing PCD and instead using only the cosine similarity between the trajectory features and the textual features as the matching score. In addition, for comparison, we emulate the approach in one-stage methods by designing two Transformer-based decoding structures. They respectively represent pre-compressing visual features into textual features and the opposite process. Here, $\mathbb{Q}$, $\mathbb{L}$ and $\mathbb{V}$ represent the query vectors, linguistic and visual features respectively.
It is evident that replacing PCD with cosine similarity results in a significant performance drop, which indicates that directly using cosine similarity to calculate matching scores in a non-CLIP semantic space can easily lead to model collapse and failure to converge. The improvement obtained on AssA by its judging all sample pairs as positive examples also proves it. Additionally, when comparing #1 and #2, it can be observed that integrating sparse visual semantics into the dense textual space achieves superior performance compared to using the visual space as the base space. However, regardless of how the order of cross-attention interactions is arranged, sequential decoding is still based on the concept of an independent modal space, and it consistently performs worse than the PCD that conducts decoding in the combined feature space.
Table 4: Impact of parameter in the loss function.
$\alpha$ $\beta$ HOTA DetA AssA
0.3 1 40.11 29.50 54.72
0.5 1 41.03 29.68 56.86
0.7 1 42.81 31.35 58.61
0.9 1 41.27 28.58 59.72
0.7 0 41.25 29.08 58.64
0.7 2 41.96 30.61 57.67
Table 5: Comparison of training and inference time. All experiments are conducted on Refer-KITTI-V2 [5]. “": The overhead when training TempRMOT only for the detection task. “††" : iKUN [8] trains at the trajectory level.
Method Epoch Train Inference HOTA
TempRMOT [5] 50 62 h 299 min 35.04
TempRMOT [5] 10 12.4 h 2.99 min -
iKUN†† [8] 100 2.51 h 115 min 10.32
JustHook 20 2.17 h 27 min 42.82
Loss Parameter
In Table 5, we evaluate the effects of different values of $\alpha$ and $\beta$, which are used to balance recall and precision and control the weighting strength of hard samples, respectively. $\alpha$ performs better when it is greater than 0.5, but when it is too high, the performance starts to degrade. Therefore, we select 0.7 as the optimal value for $\alpha$. When $\alpha$ is fixed at 0.7, we vary the value of $\beta$. It can be observed that the weight for hard samples should neither be absent ($\beta=0$) nor excessive. The highest HOTA is achieved when $\beta$ is set to 1.
Training and inference time
@@
Table 5 shows a comparison of training and inference time on our machine, where the combination of JustHook is RoBERTa [35] + ROPE Swin-T [34]. Both iKUN and TempRMOT follow the official configuration. Since iKUN only samples 4 frames of each trajectory for training in each epoch, the total epochs for iKUN are significantly high. When training TempRMOT for detection, as the dataset has only two "expressions" (car and pedestrian), TempRMOT can converge within fewer than 10 epochs. When using other detectors such as Deformable-DETR [42], the speed is generally faster. Considering that the overhead of most TBD trackers is very low, even when adding the time for training the first stage, JustHook still demonstrates a significant efficiency advantage.
More ablation experiments can be found in E
5 Conclusion
@@
In this work, we address the limitations of existing two-stage RMOT methods by proposing JustHook, a flexible and efficient hook-like framework. JustHook introduces two novel components, VFH and PCD, which enable context-aware feature aggregation and parallel decoding in a combined space, achieving significant gains in accuracy and computational efficiency. By functioning like a hook, JustHook efficiently leverages the inherent contextual aggregation capabilities of pre-trained models and introduces a novel encoder-decoder architecture for the two-stage RMOT, demonstrating strong adaptability to various emerging visual/text encoders and ensuring future-proofing.
Appendix A Dataset Details
A.1 Refer-KITTI
@@ A.2 Refer-KITTI-V2
A.3 Correction for Refer-KITTI
Refer-KITTI has suffered from multiple annotation errors for a long time. Many critical issues such as frame misalignment have been addressed in iKUN [8] and Refer-KITTI-V2 [5]. However, the problem of incomplete reference persists in the current version. For instance, in frames 231-232 of video “0002", two silver vehicles are referenced under the label “left cars in light color" but are missing in the “left cars in silver". In addition to these general errors, there are also some special errors in Refer-KITTI, and the impacts they have on one-stage methods and two-stage methods are inherently inconsistent. That is, numerous objects originally annotated in the tracking annotations of Refer-KITTI are entirely omitted in referring annotations of Refer-KITTI, such as all pedestrians in video “0001". Because one-stage methods use all tracking annotations for training while two-stage methods are only based on referring annotations, this issue leads to particularly significant adverse effects on two-stage methods.
@@ A.3 Correction for Refer-KITTI
Appendix B Limitations
Appendix C Discussion and Performance of other Trackers
Table 7: Performance comparison of multi-object tracking on KITTI [40] (using the same split as Refer-KITTI-v2 [5]). “" and “††" denote the tracker and detector versions of TempRMOT [5], respectively.
@@ Appendix C Discussion and Performance of other Trackers
As a two-stage RMOT method, the proposed JustHook can be coupled with arbitrary trackers, achieving seamless integration with any tracker, any visual backbone, and any text encoder. In order to better illustrate that the gains generated by JustHook compared with the current SOTA method TempRMOT [5] do not stem from the improvement of the tracker’s performance, in Table 7, we compare the pure multi-object tracking performance of different detector/tracker combinations (for TBD paradigm) employed in our approach with that of TempRMOT (as a pure one-stage MOT model). Here, “D-DETR" refers to DeformableDETR [42]. In the two-stage (TBD) method, TempRMOT†† serves solely as a detector and adheres to the detection paradigm, where we treat each target in each frame as a new addition. Combined with the comparison of RMOT capabilities in Table 1, it can be seen that even trackers with weaker tracking capabilities than TempRMOT can achieve performance exceeding the SOTA method TempRMOT [5] in RMOT after integrating JustHook. This demonstrates that JustHook significantly surpasses TempRMOT in matching capability, providing excellent results even when coupled with inferior trackers.
Furthermore, we compare the performance changes of JustHook on RMOT after replacing these TBD-style trackers in Table 8. JustHook achieves consistent performance improvements as the capability of the detectors / trackers increases, highlighting its flexibility and robustness.
@@
Detector & Tracker Text Encoder Visual Backbone HOTA DetA AssA LocA
@@
D-DETR [42] & OC-SORT [43] BERT [33] Swin-T [10] 37.56 27.43 51.58 89.70
Resnet34 [9] 37.83 27.95 51.36 89.92
ROPE Swin-T [34] 37.91 27.57 52.28 90.28
RoBERTa [35] Swin-T [10] 37.38 26.84 52.20 89.73
Resnet34 [9] 38.05 28.22 51.45 89.96
ROPE Swin-T [34] 38.74 29.08 51.77 89.97
D-DETR [42] & StrongSORT [44] BERT [33] Swin-T [10] 39.28 28.50 54.27 89.31
Resnet34 [9] 38.44 28.43 52.12 89.89
ROPE Swin-T [34] 38.13 28.01 52.03 90.23
RoBERTa [35] Swin-T [10] 38.95 27.73 54.83 89.32
Resnet34 [9] 38.60 28.72 52.03 89.92
ROPE Swin-T [34] 39.26 29.63 52.16 89.93
D-DETR [42] & NeuralSORT [8] BERT [33] Swin-T [10] 38.88 28.31 53.58 88.94
Resnet34 [9] 39.27 28.83 53.65 89.26
ROPE Swin-T [34] 39.27 28.61 54.07 89.58
RoBERTa [35] Swin-T [10] 38.58 27.51 54.27 88.95
Resnet34 [9] 39.63 29.38 53.61 89.24
ROPE Swin-T [34] 40.10 30.30 53.25 89.17
TempRMOT [5] & OC-SORT [43] BERT [33] Swin-T [10] 39.01 28.62 53.36 89.44
Resnet34 [9] 39.15 28.97 53.10 89.66
ROPE Swin-T [34] 39.45 29.04 53.75 90.01
RoBERTa [35] Swin-T [10] 38.32 27.49 53.58 89.45
Resnet34 [9] 39.81 29.67 53.58 89.69
ROPE Swin-T [34] 40.56 30.80 53.59 89.69
TempRMOT [5] & StrongSORT [44] BERT [33] Swin-T [10] 40.54 28.83 57.18 89.07
Resnet34 [9] 40.02 29.16 55.09 89.64
ROPE Swin-T [34] 39.98 29.16 54.98 90.01
RoBERTa [35] Swin-T [10] 39.77 27.53 57.58 89.08
Resnet34 [9] 40.58 29.83 55.34 89.69
ROPE Swin-T [34] 41.51 31.02 55.70 89.68
TempRMOT [5] & NeuralSORT [8] BERT [33] Swin-T [10] 41.14 28.78 58.94 88.97
Resnet34 [9] 41.04 29.24 57.76 89.19
ROPE Swin-T [34] 41.16 29.41 57.76 89.52
RoBERTa [35] Swin-T [10] 40.32 27.47 59.29 88.96
Resnet34 [9] 41.71 30.07 57.99 89.25
ROPE Swin-T [34] 42.81 31.35 58.61 89.22
Appendix D Broader Impact
JustHook further propels the research of RMOT, which enriches the technical means of computer vision by using language expressions as semantic cues to guide multi-object tracking predictions. This is of great significance for the development of related industries such as autonomous driving, which can improve the environmental perception ability of autonomous vehicles and enhance the safety and reliability of driving.
Appendix E More experiments
E.1 Performance comparison on Refer-Dance
We compare the performance of iKUN [8] and JustHook on Refer-Dance [8] in Table 9. Among them, JustHook uses the best combination of encoders, RoBERTa [35] and RoPE Swin Transformer-T [34]. The intermediate results of both the detector and the tracker use the versions officially provided by iKUN [8]. It can be seen that JustHook still exhibits significant performance advantages, indicating that by fully leveraging the capabilities of pre-trained models, JustHook can achieve robust improvements compared to existing methods in various scenarios.
Table 9: Performance comparison on Refer-Dance [8].
Method HOTA DetA AssA
One-stage TransRMOT [4] 9.58 4.37 20.99
Two-stage iKUN [8] 29.06 25.33 33.35
JustHook 31.83 30.31 33.47
E.2 Impact of VFH and Freezing
@@
Method Visual Backbone Text Encoder HOTA DetA AssA
JustHook +CLIP Trainable Frozen 40.76 28.93 57.57
Frozen Trainable 41.17 29.48 57.65
Frozen Frozen 40.92 29.07 57.75
Trainable Trainable 41.44 28.71 59.95
JustHook +Best Trainable Frozen 40.44 28.09 58.38
Frozen Trainable 41.05 29.73 56.81
Frozen Frozen 40.55 28.60 57.62
Trainable Trainable 42.81 31.35 58.61
In Table 10, we evaluate the impact of freezing different components in JustHook, where JustHook+CLIP represents the CLIP-R 50 [1] encoder pair, and JustHook+Best refers to the ROPE Swin-T [34] + RoBERTa [35] encoder pair. It is observed that freezing the text encoder has a more significant negative effect compared to freezing the visual backbone. This clearly demonstrates that VFH makes extremely good use of the capabilities of the pre-trained visual backbone. Thus, even when the visual backbone is frozen, there will be no significant performance loss. Moreover, VFH’s capacity to adaptively modulate contextual information facilitates a notable enhancement in performance once the visual backbone is unfrozen. However, the text encoder, whose semantic distribution diverges significantly from that of RMOT, struggles to discern subtle semantic distinctions among targets when its parameters are frozen. Therefore, training the text encoder becomes essential to enable PCD to derive a more discriminative combined feature space, thereby improving its overall discriminative power.
Additionally, freezing other encoders has a more pronounced negative impact compared to the CLIP encoders, indicating that CLIP’s inherent modality alignment effectively reduces the difficulty of constructing a unified semantic space. However, compared with the encoder pairs of CLIP, dense encoders like Swin exhibit a more pronounced performance advantage on RMOT, and training both visual / text encoders simultaneously to optimize the features of the two modalities bidirectionally consistently proves to be more beneficial.
E.3 The Scalability of JustHook
In Table 11, we evaluate the performance of JustHook with different sizes of ResNet [9] and RoPE Swin Transformer-T [34] as visual backbones. The text encoder maintains RoBERTa [35]. It can be observed that as the size of the Swin Transformer increases, the performance of JustHook does not improve significantly and even decreases. However, for convolutional networks, as the size of ResNet increases, the performance of JustHook continues to improve, but this improvement stops at ResNet34. This indicates that the scalability of visual backbones of different sizes for JustHook is not significant, but CNN-based backbones slightly outperform Transformer-based ones. This can be attributed to the limited visual information of most targets in the current Refer-KITTI-V2 dataset, which does not require a very powerful visual backbone to characterize the target’s state. When the visual backbone is too powerful, the model is more prone to overfitting. Therefore, we ultimately selected ResNet34 and ROPE Swin-T for reporting.
@@ E.3 The Scalability of JustHook
ROPE Swin [34] Tiny 42.81 31.35 58.61
Small 41.50 28.76 60.05
Base 41.16 29.43 57.70
ResNet [9] 18 40.99 28.89 58.31
34 41.71 30.07 57.99
50 40.61 27.85 59.36
E.4 The adaptability comparison between iKUN and JustHook
Visual Backbone Text Encoder HOTA DetA AssA
iKUN [8] CLIP [1] CLIP [1] 48.84 35.74 66.80
ROPE Swin [34] RoBERTa [35] 29.79 14.73 60.38
JustHook CLIP [1] CLIP [1] 52.93 42.36 66.22
ROPE Swin [34] RoBERTa [35] 53.83 43.40 66.86

### v3

@@ Introduction
We apply deep supervision to JustHook and supervise all intermediate and final outputs using Focal Loss [19], to enhance the model’s multi-scale capability and mitigate the imbalanced sample distribution in the dataset.
Experiment
Experiment Setup
Dataset and Metrics.
@@
We report the performance of JustHook on 3 datasets: Refer-Dance [7], Refer-KITTI [33] and Refer-KITTI-V2 [39]. For a fair comparison, we use Higher Order Tracking Accuracy [25] (HOTA) as the main metric, following TransRMOT [33]. HOTA comprehensively considers the detection and association capabilities of the tracking model. We also report the localization accuracy (LocA), the detection accuracy (DetA) and the association accuracy (AssA).
Implement Details.
As a two-stage RBT method, JustHook allows flexible replacement of any detector/tracker without the need for retraining. Therefore, we report only the results using the best-performing detector and tracker: TempRMOT [39] and NeuralSORT [7]. TempRMOT [39], used as the detector, is trained with detection annotations by treating each object in each frame as a new target. Structurally, it can be viewed as a Deformable-DETR [43] augmented with inter-frame interactions. NeuralSORT [7], built upon the classical SORT [1] framework, replaces the Kalman filter’s observation and transition models with neural networks, achieving excellent tracking performance. The impacts of different detectors and trackers [4; 8; 7; 43] on JustHook can be found in Supplementary Materials.
@@
In addition, JustHook can also connect with any visual backbone and text encoder. The visual backbones we used include CLIP [28], ResNet [12] and ROPE Swin [14], all of which use the pre-trained parameters on ImageNet [5], with an input size of $224\times 672$ to align as closely as possible with the aspect ratio of video. For text encoders, we test both CLIP [28] and RoBERTa [21].
During training, we iterate the trajectory segment with sample stride 2, and each sample consists of 4 frames. In neighborhood grid sampling, to simulate target loss, we sample 8 frames each time and randomly select 4 of them. For localization noise, we apply Gaussian noise with zero mean and a standard deviation equal to 0.02% of the original size to the width, height, and center of each bounding box. Random frame re-combination is performed with a probability of 1%. During inference, the sample stride is set to 1. $\hat{N}$ and $M$ are set to 36 and 10, respectively. We use AdamW [24] as the optimizer, with a learning rate of 3e-5, and train for 20 epochs. All experiments are performed on 2 NVIDIA 4090 GPUs.
Benchmark Experiments
Table 2: Performance comparison on Refer-Dance [7].
Method HOTA DetA AssA
TransRMOT [33] 9.58 4.37 20.99
iKUN [7] 29.06 25.33 33.35
JustHook-best 31.83 30.31 33.47
We compare our method with SOTA approaches on Refer-KITTI and Refer-KITTI-V2 in Table 1. As discussed in the following sections, we do not apply additional long-range enhancement to the Transformer-based backbone. JustHook achieves the best performance under various encoder configurations. Compared to the current SOTA method TempRMOT [39], our best model, JustHook-best, which adopts RoBERTa [21] as the language encoder and ROPE Swin-T [14] as the visual backbone, achieves superior performance on both Refer-KITTI and Refer-KITTI-V2. JustHook demonstrates a more significant advantage on Refer-KITTI-V2, surpassing TempRMOT by 6.9%, 7.48%, and 4.34% on HOTA, DetA, and AssA, respectively. We further report the performance of JustHook-best on Refer-Dance in Table 2, which indicates that JustHook still exhibits robust performance even in non-autonomous driving scenarios.
Table 3: Comparison of training and inference time. All experiments are conducted on Refer-KITTI-V2 [39]. “”: iKUN [7] trains at the trajectory level. The detector used is the modified TempRMOT.
Method Epoch Train Detect Infer FPS HOTA
TempRMOT [39] 50 62 h - 299 min 25.5 35.04
iKUN [7] 100 2.51 h 2.72 min 115 min 64.7 10.32
JustHook-best 20 2.17 h 2.72 min 69 min 106.3 41.94
Qualitative Results.
The qualitative results can be found in the Supplementary Materials.
Ablation Study
@@
Avg-Pool 40.69 28.14 58.97
Attention 41.36 28.96 59.20
Conv 41.51 29.42 58.71
MLP 41.94 30.45 57.92
OME.
We compare the performance of OME under Just-best setting, where Avg-Pool refers to the temporal average pooling used in iKUN [7] as a replacement for OME. As shown in Table 4, OME extracts motion information more effectively than naive pooling. However, it seems that more complex aggregation methods perform even worse than the simple MLP-based approach. This can be attributed to the fact that pre-trained visual backbones already encode rich static information within each frame. Overly complex aggregation strategies may distort the feature space, thereby hindering subsequent learning.
Table 5: Comparison of the performance about Hook and PCD on Refer-KITTI-V2.
Method Hook PCD HOTA DetA AssA
iKUN 10.32 2.17 49.77
JustHook-CLIP 35.88 22.39 57.61
JustHook-CLIP 40.14 28.26 57.16
Hook and PCD.
In Table 5, we evaluate the performance of Hook and PCD on Refer-KITTI-V2. Here, we use CLIP as encoder, and both Long-Range Enhancement and OME are disabled for all models for a fair comparison with iKUN. ✘ indicates that cosine similarity is used directly to compute the matching score instead of PCD. The results show that both PCD and Hook bring significant performance improvements. Moreover, combining with the performance of JustHook in different encoder settings, PCD can be found to perform well not only in aligned embedding spaces but also maintain strong performance in non-aligned spaces, demonstrating its potential to replace cosine similarity.
Neighboring and Long-Range Enhancement.
Table 6: Comparison of the performance about Neighboring and Long-Range Enhancement
Method $M$ HOTA DetA AssA
JustHook-CLIP 0 40.59 27.87 59.23
10 41.03 29.08 58.02
30 40.85 28.91 57.87
JustHook-R 34 0 40.97 29.85 56.39
10 41.50 29.21 59.10
30 41.39 29.14 58.92
JustHook-best 0 41.94 30.45 57.92
10 41.63 29.95 58.01
30 41.56 30.00 57.71
JustHook-best
w/o Neighbor
0 40.65 28.86 57.42
We analyze the impact of Neighboring Grid Sample and Long-Range Enhancement on Refer-KITTI-v2 in Table 6, where $M$ denotes the number of sampling points used for Long-Range Enhancement, with 0 indicating no enhancement. As observed, convolution-based visual backbones exhibit better adaptability to long-range enhancement compared to Transformer-based ones. This is because of the inherent strength of Transformers in modeling long-range dependencies, thus weakens the effect of our additional active enhancement. The number of reference points has a relatively minor impact on model performance, and setting $M=10$ is an empirical choice. In the setting without enhancement, removing the noise introduced by Hook leads to a performance drop, highlighting the importance of simulating the noise from the tracking task.
Training and inference time.
@@
Table 3 shows a comparison of efficiency on the same hardware. Both iKUN and TempRMOT follow the official configuration. Since iKUN only samples 4 frames of each trajectory for training in each epoch, its total epochs are significantly high. When training the detector, as the dataset contains only two “expressions” (car and pedestrian), our detector TempRMOT can converge in fewer than 10 epochs, instead of the 50 epochs reported in its RMOT version. The speed is generally faster in other detectors such as Deformable-DETR [43]. Considering that most TBD trackers are training-free, even when accounting for the time spent training the first stage, JustHook still demonstrates a significant efficiency advantage.
Conclusion
@@
In this work, we address the limitations of existing two-stage RBT methods by proposing JustHook, a flexible and efficient hook-like framework. JustHook introduces two novel components, Hook and PCD, which enable context-aware feature aggregation and parallel decoding in a combined space, achieving significant gains in accuracy and computational efficiency. By functioning like a hook, JustHook efficiently leverages the inherent contextual aggregation capabilities of pre-trained models and introduces a novel encoder-decoder architecture for the two-stage RMOT, demonstrating strong adaptability to various emerging visual/text encoders and ensuring future-proofing.
Dataset Details
Refer-KITTI
@@
Correction for Refer-KITTI
Refer-KITTI has suffered from multiple annotation errors for a long time. Many critical issues such as frame misalignment have been addressed in iKUN [7] and Refer-KITTI-V2 [39]. However, the problem of incomplete reference persists in the current version. For instance, in frames 231-232 of video “0002”, two silver vehicles are referenced under the label “left cars in light color” but are missing in the “left cars in silver”. In addition to these general errors, there are also some special errors in Refer-KITTI, and the impacts they have on one-stage methods and two-stage methods are inherently inconsistent. That is, numerous objects originally annotated in the tracking annotations of Refer-KITTI are entirely omitted in referring annotations of Refer-KITTI, such as all pedestrians in video “0001”. Because one-stage methods use all tracking annotations for training while two-stage methods are only based on referring annotations, this issue leads to particularly significant adverse effects on two-stage methods.
@@
Table 8: Performance comparison of multi-object tracking on KITTI [11] (using the same split as Refer-KITTI-V2 [39]). “” and “††” denote the tracker and detector versions of TempRMOT [39], respectively.
@@
Detector
& Tracker
Text
Encoder
Visual
Backbone
HOTA DetA AssA LocA
@@
D-DETR [43]
& OC-SORT [4]
RoBERTa [21] Resnet34 [12] 38.20 27.95 52.34 89.91
ROPE Swin-T [14] 37.94 28.59 50.53 89.72
CLIP [28] CLIP [28] 36.13 26.31 49.75 89.99
D-DETR [43]
& StrongSORT [8]
RoBERTa [21] Resnet34 [12] 39.70 29.01 54.49 89.53
ROPE Swin-T [14] 39.66 29.85 52.85 89.35
CLIP [28] CLIP [28] 37.44 27.33 51.45 89.65
D-DETR [43]
& NeuralSORT [7]
RoBERTa [21] Resnet34 [12] 39.62 28.86 54.55 89.18
ROPE Swin-T [14] 39.46 29.67 52.64 88.98
CLIP [28] CLIP [28] 37.76 27.23 52.49 89.34
TempRMOT [39]
& OC-SORT [4]
RoBERTa [21] Resnet34 [12] 39.57 29.01 54.13 89.71
ROPE Swin-T [14] 39.69 30.10 52.51 89.47
CLIP [28] CLIP [28] 37.47 27.36 51.46 89.85
TempRMOT [39]
& StrongSORT [8]
RoBERTa [21] Resnet34 [12] 41.03 29.29 57.61 89.33
ROPE Swin-T [14] 41.37 30.51 56.26 89.09
CLIP [28] CLIP [28] 38.89 27.90 54.34 89.49
TempRMOT [39]
& NeuralSORT [7]
RoBERTa [21] Resnet34 [12] 41.50 29.21 59.10 89.25
ROPE Swin-T [14] 41.94 30.45 57.92 89.00
CLIP [28] CLIP [28] 41.03 29.08 58.2 89.03
Discussion and Performance of other Trackers
In order to better illustrate that the gains generated by JustHook do not heavily rely on the improvement of the tracker’s performance, in Table 8, we compare the pure multi-object tracking performance of different detector/tracker combinations (for TBD paradigm) employed in our approach with that of TempRMOT (as a pure one-stage MOT model). Here, “D-DETR” refers to DeformableDETR [43]. In the two-stage (TBD) method, TempRMOT†† serves solely as a detector and adheres to the detection paradigm, where we treat each target in each frame as a new addition. Combined with the comparison of RMOT capabilities in paper, it can be seen that even trackers with weaker tracking capabilities than TempRMOT can achieve performance exceeding the SOTA method TempRMOT [39] in RMOT after integrating JustHook. This demonstrates that JustHook significantly surpasses TempRMOT in matching capability, providing excellent results even when coupled with inferior trackers.
Furthermore, we compare the performance changes of JustHook on RMOT after replacing these TBD-style trackers in Table 9. The version of JustHook is the same as that in the paper. JustHook achieves consistent performance improvements as the capability of the detectors/trackers increases, highlighting its flexibility and robustness.
More experiments
Visual
Backbone
Text
Encoder
HOTA DetA AssA
iKUN [7] CLIP [28] CLIP [28] 48.84 35.74 66.80
ROPE Swin [14] RoBERTa [21] 29.79 14.73 60.38
JustHook CLIP [28] CLIP [28] 53.49 43.62 65.68
ROPE Swin [14] RoBERTa [21] 53.91 44.02 66.09
Impact of Freezing
In Table 11, we evaluate the impact of freezing different components in JustHook, where JustHook-CLIP represents the CLIP-R 50 [28] encoder pair, and JustHook-best refers to the ROPE Swin-T [14] + RoBERTa [21] encoder pair.
It can be observed that different encoders respond differently to parameter freezing. For the CLIP pre-trained encoder, which is naturally aligned across modalities, training only one side disrupts the original embedding space, leading to performance even worse than freezing both sides entirely. In contrast, for RoBERTa and ROPE Swin-T, freezing the text encoder results in a significantly larger performance drop compared to freezing the visual encoder. This can be attributed to the fact that the textual distribution in the Refer-KITTI-V2 dataset differs more from RoBERTa’s pre-training corpus, being focused solely on the autonomous driving, than the visual gap does. Therefore, the model requires more learning on the text side to adapt to this domain shift.
Moreover, Table 11 also shows that JustHook generally outperforms CLIP when using denser visual backbones. However, the efficiency gain brought by freezing the CLIP encoder is unmatched by any of the other encoders. Therefore, by flexibly selecting different encoders based on the actual hardware environment, JustHook demonstrates strong potential for broad real-world applicability.
@@
Method
Visual
Backbone
Text
Encoder
HOTA DetA AssA
JustHook
-CLIP
Trainable Frozen 40.23 28.96 56.05
Frozen Trainable 40.08 28.44 56.60
Frozen Frozen 40.70 29.03 57.22
Trainable Trainable 41.03 29.08 58.02
JustHook
-best
Trainable Frozen 40.79 28.34 58.88
Frozen Trainable 41.59 29.78 58.20
Frozen Frozen 40.62 27.81 59.46
Trainable Trainable 41.94 30.45 57.92
The Scalability of JustHook
In Table 12, we evaluate the performance of JustHook with different sizes of ResNet [12] and ROPE Swin Transformer-T [14] as visual backbones. The text encoder maintains RoBERTa [21]. It can be observed that as the size of the Swin Transformer increases, the performance of JustHook does not improve significantly and even decreases. However, for convolutional networks, as the size of ResNet increases, the performance of JustHook also improves, but this improvement stops at ResNet34. This indicates that the scalability of visual backbones of different sizes for JustHook is not significant, but CNN-based backbones slightly outperform Transformer-based ones. This can be attributed to the limited visual information of most targets in the current Refer-KITTI-V2 dataset, which does not require a very powerful visual backbone to characterize the target’s state. When the visual backbone is too powerful, the model is more prone to overfitting. Therefore, we finally selected ResNet34 and ROPE Swin-T for reporting.
@@
ROPE Swin [14] Tiny 41.94 30.45 57.92
Small 41.27 30.02 56.89
Base 41.20 28.97 58.72
ResNet [12] 18 40.90 28.80 58.20
34 41.50 29.21 59.10
50 41.17 29.70 57.20
The reliance on CLIP of iKUN and JustHook
@@
Limitations
Broader Impact
JustHook further accelerates the research of RMOT, which enriches the technical means of computer vision by using language expressions as semantic cues to guide multi-object tracking predictions. This is of great significance for the development of related industries such as autonomous driving, which can improve the environmental perception ability of autonomous vehicles and enhance the safety and reliability of driving.
Qualitative Results